OpenStudy (anonymous):

what is equivalent to 63y-4/14y^2

OpenStudy (anonymous):

\[\frac{ 63y-4 }{ 14y^2 }\]

OpenStudy (anonymous):

um what are u trying to get?, that expression can be represented in many ways

OpenStudy (anonymous):

Our goal here is to divide both of the two terms on top by the one term on the bottom, so: \[\frac{ 63y }{ 14y^2 } - \frac{ 4 }{ 14y^2 }\] Simplify and we'll get: \[\frac{ 63 }{ 14y } - \frac{ 2 }{ 7y^2 }\]